OpenStudy (anonymous):

Find the slope of the line that goes through the given points (-6,7) and (-3,-5)

OpenStudy (anonymous):

y2-y1/x2-x1 = -5 - 7/-3--6 = -12/3 = -4

OpenStudy (anonymous):

Supposed to type an integer or a simple fraction?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):